OpenStudy (anonymous):

Solve for x , in the interval 0

OpenStudy (anonymous):

sin x + cos x = sin x + sin (Pi/2 - x) , because cos x = sin (pi/2 - x) Next, use the identity: sin a + sin b = 2sin (a + b)/2*cos (a - b)/2. sin x + sin (Pi/2 - x) = 2*sin (Pi/4)* sin (x + Pi/4) = ......

OpenStudy (anonymous):

Your equation corresponds to the intersections of the line x + y = 0 and the unit circle. You get one solution in the second and fourth quadrants.  The angles are 135 degrees and 315 degrees.